# MathGenie: Generating Synthetic Data with Question Back-translation for Enhancing Mathematical Reasoning of LLMs

### v1

@@ Abstract
$*{}^{*}$qual contribution ${}^{\dagger}$orresponding author
@@ 3 Experiments
ChatGPT-3.5
80.8
35.5
83.0
GPT-4
92.0
42.5
97.0
GPT-4 Code
97.0
69.7
PaLM-2
80.7
34.3
@@
Mammoth
CodeLlama
7B
59.4
33.4
71.4
45.9
55.4
53.1
MathCoder
CodeLlama
7B
67.8
30.2
70.7
49.6
55.8
54.8
ToRA
CodeLlama
7B
72.6
44.6
70.4
36.0
68.1
58.3
MathGenieLM
CodeLlama
7B
71.5
39.7
80.2
69.1
69.5
66.0
Mammoth
Llama-2
7B
53.6
31.5
67.7
41.2
46.3
48.1
MathCoder
Llama-2
7B
64.2
23.3
71.5
47.5
46.9
50.7
ToRA
Llama-2
7B
68.8
40.1
68.2
29.2
58.3
52.9
MathGenieLM
Llama-2
7B
71.7
33.0
78.5
61.4
65.0
61.9
MathGenieLM
Llemma
7B
76.0 48.3 81.3 85.0 76.6 73.4
MathGenieLM
Mistral
7B
80.5 45.1 83.3 79.4 71.8 72.0
Mammoth
CodeLlama
13B
64.7
36.3
73.7
47.1
61.5
56.7
MathCoder
CodeLlama
13B
74.1
35.9
78.0
60.7
62.5
62.2
ToRA
CodeLlama
13B
75.8
48.1
75.7
37.9
70.3
61.6
MathGenieLM
CodeLlama
13B
78.5
40.3
84.5 76.7 65.7 69.1
Mammoth
Llama-2
13B
62.0
34.2
72.4
43.2
49.2
52.2
MathCoder
Llama-2
13B
72.6
29.9
76.9
62.3
54.7
59.3
ToRA
Llama-2
13B
72.7
43.0
72.9
45.7
62.6
59.4
MathGenieLM
Llama-2
13B
80.4 43.8 83.5 78.4 72.7 71.8
Mammoth
CodeLlama
34B
72.7
43.6
84.3
51.8
65.4
63.6
MathCoder
CodeLlama
34B
81.7
45.2
82.5
65.8
75.9
70.2
ToRA
CodeLlama
34B
80.7
50.8
80.5
50.2
77.9
68.0
MathGenieLM
CodeLlama
34B
86.2 51.4 86.9 85.8 78.4 77.7
MathGenieLM
Llemma
34B
84.1 55.1 87.4 89.3 82.9 79.8
WizardMath
Llama-2
70B
81.6
22.7
Mammoth
Llama-2
70B
76.9
41.8
82.4
51.4
55.6
61.6
MathCoder
Llama-2
70B
83.9
45.1
84.9 77.0 74.4 73.1
ToRA
Llama-2
70B
84.3 49.7
82.7
73.3
72.6
72.5
MathGenieLM
Llama-2
70B
88.4 51.2 87.7 89.1 76.0 78.5
MathGenieLM
Mixtral
8x7B
87.0
53.7
88.9
89.9
81.7
80.2
MathGenieLM
InternLM2
20B
87.7
55.7
87.3
88.5
85.1
80.9
@@
Model Base Voting
$k$-path
In-Domain Out-of-Domain
@@
ToRA
Llama-2 70B
50
88.3
56.9
MathGenieLM Llama-2 70B
88.4
51.2
87.7
89.1
76.0
78.5
MathGenieLM Llama-2 70B
10
91.7(+3.3)
63.3(+12.1)
94.0(+6.3)
95.9(+6.8)
87.2(+11.2)
86.4(+7.9)
@@ 3.3 Ablation Study
w/ verification filtering
79.3
43.8
61.6
w/o verificatioin filtering
80.5(+1.2)
45.1(+1.3)
62.8(+1.2)
@@
MetaMath
79.5
44.6
78.4
79.6
67.9
70.0
Direct question augmentation (w/o sol.)
78.8
43.0
84.0
77.0
70.2
70.6
Direct question augmentation (w/ sol.)
79.2
44.2
83.6
72.0
68.6
69.5
@@ 3.4 Accuracy of Verified Inference
Verify (twice) Accuracy
88.6
55.8
88.7
91.2
81.1
81.1
@@
Voting (3-path) Accuracy
88.6
53.8
92.0
90.9
81.5
81.4
@@ Appendix C Prompt of Verification
**Question**:
{question}
**Solution**:
{solution}
@@ Appendix D Analysis of Code-integrated Verification
Accuracy
86.4
49.5
86.0
88.1
73.6
76.7
Precision
89.3(+2.9)
64.5(+15.0)
88.3(+2.3)
90.8(+2.7)
79.9(+3.2)
82.6(+5.9)
Recall
98.5
94.7
97.4
98.2
97.1
97.2

### v2

@@ Abstract
Equal contribution Corresponding author
@@ 3 Experiments
ChatGPT-3.5 - - 80.8 35.5 83.0 - - -
GPT-4 - - 92.0 42.5 97.0 - - -
GPT-4 Code - - 97.0 69.7 - - - -
PaLM-2 - - 80.7 34.3 - - - -
@@
Mammoth CodeLlama 7B 59.4 33.4 71.4 45.9 55.4 53.1
MathCoder CodeLlama 7B 67.8 30.2 70.7 49.6 55.8 54.8
ToRA CodeLlama 7B 72.6 44.6 70.4 36.0 68.1 58.3
MathGenieLM CodeLlama 7B 71.5 39.7 80.2 69.1 69.5 66.0
Mammoth Llama-2 7B 53.6 31.5 67.7 41.2 46.3 48.1
MathCoder Llama-2 7B 64.2 23.3 71.5 47.5 46.9 50.7
ToRA Llama-2 7B 68.8 40.1 68.2 29.2 58.3 52.9
MathGenieLM Llama-2 7B 71.7 33.0 78.5 61.4 65.0 61.9
MathGenieLM Llemma 7B 76.0 48.3 81.3 85.0 76.6 73.4
MathGenieLM Mistral 7B 80.5 45.1 83.3 79.4 71.8 72.0
Mammoth CodeLlama 13B 64.7 36.3 73.7 47.1 61.5 56.7
MathCoder CodeLlama 13B 74.1 35.9 78.0 60.7 62.5 62.2
ToRA CodeLlama 13B 75.8 48.1 75.7 37.9 70.3 61.6
MathGenieLM CodeLlama 13B 78.5 40.3 84.5 76.7 65.7 69.1
Mammoth Llama-2 13B 62.0 34.2 72.4 43.2 49.2 52.2
MathCoder Llama-2 13B 72.6 29.9 76.9 62.3 54.7 59.3
ToRA Llama-2 13B 72.7 43.0 72.9 45.7 62.6 59.4
MathGenieLM Llama-2 13B 80.4 43.8 83.5 78.4 72.7 71.8
Mammoth CodeLlama 34B 72.7 43.6 84.3 51.8 65.4 63.6
MathCoder CodeLlama 34B 81.7 45.2 82.5 65.8 75.9 70.2
ToRA CodeLlama 34B 80.7 50.8 80.5 50.2 77.9 68.0
MathGenieLM CodeLlama 34B 86.2 51.4 86.9 85.8 78.4 77.7
MathGenieLM Llemma 34B 84.1 55.1 87.4 89.3 82.9 79.8
WizardMath Llama-2 70B 81.6 22.7 - - - -
Mammoth Llama-2 70B 76.9 41.8 82.4 51.4 55.6 61.6
MathCoder Llama-2 70B 83.9 45.1 84.9 77.0 74.4 73.1
ToRA Llama-2 70B 84.3 49.7 82.7 73.3 72.6 72.5
MathGenieLM Llama-2 70B 88.4 51.2 87.7 89.1 76.0 78.5
MathGenieLM Mixtral 8x7B 87.0 53.7 88.9 89.9 81.7 80.2
MathGenieLM InternLM2 20B 87.7 55.7 87.3 88.5 85.1 80.9
@@
Model Base Voting $k$-path In-Domain Out-of-Domain
@@
ToRA Llama-2 70B 50 88.3 56.9 - - - -
MathGenieLM Llama-2 70B - 88.4 51.2 87.7 89.1 76.0 78.5
MathGenieLM Llama-2 70B 10 91.7(+3.3) 63.3(+12.1) 94.0(+6.3) 95.9(+6.8) 87.2(+11.2) 86.4(+7.9)
@@ 3.3 Ablation Study
w/o verification filtering 79.3 43.8 61.6
w/ verificatioin filtering 80.5(+1.2) 45.1(+1.3) 62.8(+1.2)
@@
MetaMath 79.5 44.6 78.4 79.6 67.9 70.0
Direct question augmentation (w/o sol.) 78.8 43.0 84.0 77.0 70.2 70.6
Direct question augmentation (w/ sol.) 79.2 44.2 83.6 72.0 68.6 69.5
@@ 3.4 Accuracy of Verified Inference
Verify (twice) Accuracy 88.6 55.8 88.7 91.2 81.1 81.1
@@
Voting (3-path) Accuracy 88.6 53.8 92.0 90.9 81.5 81.4
@@ 5 Conclusion
Acknowledgement
This project is funded in part by National Key R&D Program of China Project 2022ZD0161100, by the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK, by General Research Fund of Hong Kong RGC Project 14204021. Hongsheng Li is a PI of CPII under the InnoHK.
@@ Appendix C Prompt of Verification
**Question**: {question} **Solution**: {solution} Above is a math problem and its solution. Please use code to verify the solution above.
@@ Appendix D Analysis of Code-integrated Verification
Accuracy 86.4 49.5 86.0 88.1 73.6 76.7
Precision 89.3(+2.9) 64.5(+15.0) 88.3(+2.3) 90.8(+2.7) 79.9(+3.2) 82.6(+5.9)
Recall 98.5 94.7 97.4 98.2 97.1 97.2